A die is rolled 6 times. Find the probability of rolling no more than 3 three's.
Answer #1

This is a binomial distribution question with
n = 6
p = 0.1667     (probability of getting a three = 1/6)
q = 1 - p = 0.8333
where
P(X = 1) =
\\ a)P(X \le 3), n = 6, p = 0.1667 \\ P(X \le 3) = P(0 \le X \le 3) \\ P(X \le 3) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) \\ P(X \le 3) = \binom{6}{0} 0.1667^{0} 0.8333^{6-0} + \binom{6}{1} 0.1667^{1} 0.8333^{6-1} + \binom{6}{2} 0.1667^{2} 0.8333^{6-2} + \binom{6}{3} 0.1667^{3} 0.8333^{6-3} \\ P(X \le 3) = 1 (0.1667^{0}) (0.8333^{6}) + 6.0 (0.1667^{1}) (0.8333^{5}) + 15.0 (0.1667^{2}) (0.8333^{4}) + 20.0 (0.1667^{3}) (0.8333^{3}) \\ P(X \le 3) = 1 (1.0) (0.3348) + 6.0 (0.1667) (0.4018) + 15.0 (0.0278) (0.4822) + 20.0 (0.0046) (0.5786) \\ P(X \le 3) = \textbf{0.3348176092} + \textbf{0.4018775624} + \textbf{0.2009870084} + \textbf{0.0536093993} \\ P(X \le 3) = \textbf{0.9913} \\ \par\noindent\rule{\textwidth}{0.4pt}

Please hit thumbs up if the answer helped you